### v2

# WikiBigEdit: Understanding the Limits of Lifelong Knowledge Editing in LLMs

Abstract
Keeping large language models factually up-to-date is crucial for deployment, yet costly retraining remains a challenge. Knowledge editing offers a promising alternative, but methods are only tested on small-scale or synthetic edit benchmarks. In this work, we aim to bridge research into lifelong knowledge editing to real-world edits at a practically relevant scale. We first introduce WikiBigEdit; a large-scale benchmark of real-world Wikidata edits, built to automatically extend lifelong for future-proof benchmarking. In its first instance, it includes over 500K question-answer pairs for knowledge editing alongside a comprehensive evaluation pipeline. Finally, we use WikiBigEdit to study existing knowledge editing techniques’ ability to incorporate large volumes of real-world facts and contrast their capabilities to generic modification techniques such as retrieval augmentation and continual finetuning to acquire a complete picture of the practical extent of current lifelong knowledge editing. [^1] Code available at https://github.com/ExplainableML/WikiBigEdit.
Machine Learning, ICML
1 Introduction
Keeping large language models factually up-to-date is a critical challenge in their practical deployment (Hartvigsen et al., 2023; Meng et al., 2023, 2022; Wang et al., 2024a; Zhang et al., 2024) due to given knowledge cutoffs and limited tail-end concept coverage (Cheng et al., 2024; Zheng et al., 2024; Zhang et al., 2024; Udandarao et al., 2024) incurring inaccuracies. Finding ways to inject or update new and revised facts is crucial; Particularly for factuality-critical applications like medicine (Bao et al., 2023; Jeong et al., 2024), law (Colombo et al., 2024; Lai et al., 2024) or education (Team et al., 2024c; Wang et al., 2024b).
However, frequent full model retraining is prohibitively expensive. To address this, lifelong knowledge editing techniques (Hartvigsen et al., 2023; Meng et al., 2023; Wang et al., 2024a) have recently emerged. In controlled benchmark studies, they have shown success incorporating new facts along three key metrics: (1) generalization beyond merely memorizing question-answer pairs, (2) locality of edits to maintain pretraining knowledge, and (3) retention of all previously seen edits.
Unfortunately, existing benchmarks are small [^2] CounterFact (Meng et al., 2023) at 20K, ZsRE (Levy et al., 2017) at 1K, SelfCheckGPT (Manakul et al., 2023) at 600., synthetic, and precede the knowledge cutoff of modern LLMs. While more recent efforts Jang et al. (2022); Khodja et al. (2024) avoid synthetic edits by using wikidata, they remain small-scale (around 20K), and have yet to examine the difficulty of edits, limiting our understanding of really lifelong knowledge editing capabilities at practically-needed scales. LLMs are nowadays trained on trillions of text tokens (e.g. Llama-3 (et al., 2024a) on 15T), and studying lifelong factuality updates demands large volumes of recent real-world knowledge edits.
To better understand lifelong knowledge editing at a scale and scenario of practical importance, we introduce WikiBigEdit - a large-scale, lifelong benchmark (Prabhu et al., 2024; Ghosh et al., 2024) with a fully automated dataset extraction pipeline that continuously expands with new real-world factual edits to ensure future-proof factuality probing. At its inception, WikiBigEdit covers over 500K high-quality question-answer pairs (around 7 million tokens; magnitudes larger than standard edit benchmarks) spanning eight time-intervals over five months (February - July 2024) - reflecting a typical interval between subsequent version releases of large-scale LLMs [^3] E.g. Llama 1 (Touvron et al., 2023) to LLama 2 (et al., 2023), DeepSeek-V1 (et al., 2024b) to V2 (et al., 2024c) and DeepSeek-V3 (et al., 2024d). to better emulate the deployment time required to cover with factual updates.
Derived from periodic changes to Wikidata knowledge graphs (Jang et al., 2022; Khodja et al., 2024), WikiBigEdit covers a large range of factual edits and refinements. Moreover, WikiBigEdit introduces comprehensive evaluation axes going beyond standard knowledge editing performance measures, incorporating locality checks, multi-hop reasoning across incorporated edits as well as standard and complex generalization tests. We use WikiBigEdit to explore the properties of real-world knowledge edits and conduct an extensive study of knowledge editing techniques for WikiBigEdit-scale updates, focusing on their ability to integrate numerous sequential edits while maintaining previous knowledge and factual accuracy beyond simple memorization. Additionally, we compare these methods against other generic modification techniques such as retrieval augmentation (Vu et al., 2023; Prabhu et al., 2023a; Gui et al., 2024) and continual finetuning (Ibrahim et al., 2024; Roth et al., 2024). Our study reveals limited transfer of knowledge editing techniques to practically scalable lifelong updates and highlights that even when evaluated around knowledge editing desiderata, standard approaches such as retrieval augmentation or continual finetuning with model merging can perform better.
Overall, we make the following contributions to the study of real-world lifelong knowledge editing as scale:
  • We present WikiBigEdit: a large-scale benchmark of 500k high-quality question-answer pairs spanning eight timesteps over six months, derived from real-world factual changes in Wikidata to both analyze and simulate real-world knowledge edit tasks at a practically relevant scale.
  • A fully automated extraction pipeline which continuously extracts new, suitable factual edits from Wikidata knowledge graphs, enabling the benchmark to evolve with new knowledge updates continuously.
  • Using WikiBigEdit, we thoroughly analyze the capability of existing lifelong knowledge editing methods to conduct lifelong edits at scale; contrasted against retrieval augmentation and continual finetuning to understand limits in relation to other established approaches.
2 Related Works
Benchmark Size Date Data Source Task Lifelong Mhop
NQ 307K 2016 Google Search queries Open factual QA
Trivia QA 650K Various Trivia sources (web) Trivia QA
MS MARCO 1M+ 2016 Bing Search queries Search queries
Hotpot QA 112K 2018 Wikipedia (curated) Multi-hop QA
FEVER 185K 2018 Human-written claims Fact verification
EX-FEVER 60K 2023 Hyperlinked Wikipedia Mhop fact verif.
ELI5 270K 2019 Reddit (ELI5 subreddit) Long-form QA
WikiQA 3.5K 2015 Wikipedia Factoid QA
DatedData 200K Various Varied web sources Temporal QA
StreamingQA 150K 2007-2020 WMT news articles Event-based QA
ArchivalQA 530K 1985-2008 Historical news archives Fact-based QA
Hello Fresh 30K 2023-2024 X and Wikipedia Fact verification
CLARK 1.4K 2021-2024 Wikipedia Knowledge QA
PopQA 14k 2023 Wikipedia, Wikidata Fact-based QA
TemporalWiki 7K 2021 Wikipedia, Wikidata Temporal QA
WikiFactDiff 20k 2021-2023 Wikidata Factual cloze tests
WikiBigEdit 502K 2024 Wikidata Fact-based QA
Table 1: Comparison of QA and edit benchmarks in terms of size, date, question types, sources, and lifelong applicability; highlighting the unique placement of our WikiBigEdit benchmark.
(Lifelong) Knowledge Editing Benchmarks. Traditional QA datasets like NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and MS MARCO (Bajaj et al., 2018) target open-domain factual QA, HotpotQA (Yang et al., 2018) and EX-FEVER (Ma et al., 2024) assess multi-hop reasoning. Fan et al. (2019); Li et al. (2024); Cheng et al. (2024); Mallen et al. (2023) focus on long-form explanatory questions and time-sensitive QA. StreamingQA (Liška et al., 2022), ArchivalQA (Wang et al., 2021) or HelloFresh (Franzmeyer et al., 2024) address evolving or time-sensitive, event-driven data, while TiC-LM (Li et al., 2025) explores continual pretraining of LLMs. Knowledge editing benchmarks assess the ability to integrate new facts without retraining. ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023) focus on synthetic updates, while AToKE (Yin et al., 2024) and ChronoEdit (Ge et al., 2024) evaluate temporal consistency but remain limited in scale and scope. ELKEN (Peng et al., 2024) and EVEDIT (Liu et al., 2024b) generate edits based on real-world events. TemporalWiki (Jang et al., 2022) and WikiFactDiff [^4] While WikiFactDiff contains over 450K triple–text pairs, only a 20K subset ($\text{WFD}_{\text{repl}}$) is commonly used for editing due to its defined evaluation protocol. (Khodja et al., 2024) are more closely aligned with WikiBigEdit. Both evaluate temporal update consistency using consecutive Wikidata snapshots. WikiBigEdit provides over a magnitude more samples (Table 2), future-proofing and a comprehensive evaluation along all editing desiderata.
Knowledge Editing. Approaches for integrating new knowledge into LLMs include local modification methods such as ROME (Meng et al., 2023) and MEMIT (Meng et al., 2022), global optimization techniques like MEND (Mitchell et al., 2021), and external memorization strategies such as SERAC (Mitchell et al., 2022), GRACE (Hartvigsen et al., 2023), RECIPE (Chen et al., 2024), and WISE (Wang et al., 2024a). While these techniques support efficient model updates, they often suffer from catastrophic forgetting and knowledge degradation when applied to large-scale or sequential edits (Hsueh et al., 2024; Gupta et al., 2024b; Yang et al., 2024b). Our benchmark offers a practical framework to examine these limitations under realistic, evolving knowledge conditions.
Retrieval Augmented Generation (RAG). By combining a retrieval module with a generative language model, this class of methods addresses issues such as outdated knowledge and untraceable generations (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020). RAG approaches have shown strong performance on knowledge-intensive tasks, even at smaller model scales (Borgeaud et al., 2022; Izacard et al., 2022), and enable attribution by linking generated outputs to retrieved sources (Rashkin et al., 2022; Honovich et al., 2022). Moreover, RAG supports dynamic knowledge updates, mitigating context-memory conflicts when parametric information becomes stale (Vu et al., 2023). However, RAG faces limitations in multi-hop reasoning and compositional retrieval. More advanced variants - such as HippoRAG (Gutiérrez et al., 2025) or agent-based RAG systems - aim to address these limitations by improving reasoning over retrieved information.
Continual Finetuning and Model Merging. Continual fine-tuning has emerged as a reliable approach for general large-scale model updates (Garg et al., 2024; Prabhu et al., 2023b; Ibrahim et al., 2024). Roth et al. (2024) show that parameter-efficient finetuning via e.g. LoRA (Hu et al., 2021) or GaLore (Zhao et al., 2024), can be very effective, leveraging regularization provided by limiting trainable parameter counts (Thede et al., 2024). Parameter-selective techniques, including BitFit (Zaken et al., 2022) and LNFit (Min et al., 2023), further reduce computational overhead by targeting specific parameter subsets. Additionally, regularization strategies from continual learning, such as EWC (Kirkpatrick et al., 2017) and SI (Zenke et al., 2017), have shown notable success in mitigating forgetting during the continual adaptation of foundation models. Model merging (Wortsman et al., 2022; Ilharco et al., 2022; Ramé et al., 2024) has shown promising gains coupled with finetuning for continual tasks (Stojanovski et al., 2022; Marouf et al., 2024; Kozal et al., 2024; Marczak et al., 2024; Dziadzio et al., 2024). We investigate and contrast these approaches against knowledge editing at scale.
3 WikiBigEdit
Figure 1: Automated Dataset Generation Pipeline. (1) Comparing two Wikidata knowledge graph snapshots; extract sets of unchanged and changed subject-relation-object triplets, (2) filter for quality, (3) generate locality probes, (4) produce multi-hop reasoning qintuples, (5) convert into question-answer pairs using e.g. GPT-3.5, (6) introduce rephrased and persona-based evaluations to further test generalization, conduct (7) final quality assurance. Put together, this can continuously extend WikiBigEdit with additional time-specific high-quality knowledge edit data.
WikiBigEdit addresses three key aspects: (1) it extracts a large number of realistic knowledge edits, (2) incorporates extensive evaluations to assess generalization, reasoning, and locality, and (3) operates fully automatically. The complete pipeline is shown in Figure 1.
Problem Definition.
The task of large-scale, lifelong knowledge editing requires updating a pre-trained LLM $f^{0}$ with a stream of factual updates $u_{1},u_{2},\ldots,u_{T}]$, where each update $u_{t}=(q_{t},a_{t})$ consists of a question $q_{t}$ and its factual answer $a_{t}$. These updates are grouped into $B$ sequential batches $U_{1},U_{2},\ldots,U_{B}]$, where each batch can encompass anything from a single edit to multiple. For each batch update $b$ (also denoted as timestep in this work), the model $f^{b-1}$, trained on updates from prior batches $U_{<b}$, is further updated with the current batch $U_{b}$ to produce $f^{b}$. The model must meet three key objectives: (1) integrate updates from $U_{b}$, (2) retain knowledge from $U_{<b}$, and (3) ensure that the semantic understanding required to correctly solve previously seen edits generalizes beyond simple memorization of exact question-answer pairings.
3.1 Lifelong Automated Extraction of Factual Edits
As a proxy for world knowledge, we leverage Wikidata knowledge graphs, following recent works (Jang et al., 2022; Khodja et al., 2024). By performing pairwise comparisons of snapshots taken at different points in time, we identify changes that approximate real-world knowledge evolution. We thus automatically generate WikiBigEdit using our proposed seven-step pipeline, as follows:
(1) Periodic Snapshot Acquisition. The automated pipeline is initiated through periodic downloads of most recent snapshots of the Wikidata knowledge graph, and paired with the last downloaded snapshot. Pairwise comparison of these snapshots allows us to extract two distinct sets of subject-relation-object triplets $\mathcal{T}=(\text{subj},\text{rel},\text{obj})$, $\mathcal{S}=\{\mathcal{T}_{i}\}$: $\mathcal{S}_{\text{static}}$ representing unchanged facts and $\mathcal{S}_{\text{changed}}$ capturing updated or newly added information (c.f. top of Figure 1). While $\mathcal{S}_{\text{changed}}$ will be used to create the knowledge edits to be incorporated into a base model, $\mathcal{S}_{\text{fixed}}$ is primarily utilized for evaluations. Each triplet describes a given relation (such as “capital”, “manufacturer”, …) between a subject (e.g. “Norway”) and object (s.a. “Oslo”). A detailed overview is provided in the supplementary.
(2) Initial Filtering. To ensure high-quality triplets, we apply several filtering steps to both $\mathcal{S}_{\text{static}}$ and $\mathcal{S}_{\text{changed}}$: (i) We exclude triplets with circular dependencies, (ii) where the subject or object entities consist of non-Roman characters, (iii) single characters, (iv) overly long phrases (e.g., more than five words) and (v) where a single subject-relation combination maps to multiple objects to avoid ambiguity.
(3) Generation of Locality Probes. The filtered $\mathcal{S}_{\text{changed}}$ defines the knowledge updates. To evaluate whether knowledge edits remain local, i.e. do not influence (related) existing knowledge, we derive a locality set $\mathcal{S}_{\text{locality}}$ from $\mathcal{S}_{\text{static}}$ and $\mathcal{S}_{\text{changed}}$; where each factual triplet in the changed set is paired with a triplet semantically similar in subject and relation but with a distinct object. For each factual update, this gives a corresponding question whose answer should remain unaffected to probe the locality of knowledge edits.
(4) Inclusion of Multi-Hop Reasoning Tuples. To enable reasoning evaluation, we further extract multi-hop (mhop) tuples $\mathcal{S}_{\text{mhop}}$ from $\mathcal{S}_{\text{changed}}$. In particular, we identify pairs of factual triplets $(\mathcal{T}_{i},\mathcal{T}_{j})$ where $\mathcal{T}_{i}^{\text{obj}}$ serves as the subject of the second triplet, $\mathcal{T}_{j}^{\text{subj}}$; for example $\mathcal{T}_{i}=$ (podcast, named after, iPod) and $\mathcal{T}_{j}=$ (iPod, manufacturer, Apple). The triplets $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are then concatenated to form two-hop factual qintuples $\mathcal{T}_{ij}$, e.g. (podcast, named after, iPod, manufacturer, Apple). By restricting the extraction of $\mathcal{S}_{\text{mhop}}$ to $\mathcal{S}_{\text{changed}}$, we ensure a fair evaluation, as the model will have been updated with both individual facts prior to the reasoning task. We find that for a filtered batch $\mathcal{S}_{\text{changed}}$, we can recombine around $3.9\%$ update triplets into $\mathcal{S}_{\text{mhop}}$.
(5) Generation of Question-Answer Edit Pairs. For $\mathcal{S}_{\text{changed}}$, $\mathcal{S}_{\text{locality}}$, and $\mathcal{S}_{\text{mhop}}$, we generate corresponding question-answer pairs using GPT-4o-mini (Brown et al., 2020) (GPT-3.5 for update and locality) through few-shot prompting (for all prompt templates, please see the supplementary). For $\mathcal{S}_{\text{mhop}}$, we generate questions that omit the middle entity of the two-hop factual quintuple. Using this, we generate the corresponding question-answer sets $\mathcal{S}^{\text{QA}}_{\text{changed}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$. $\mathcal{S}^{\text{QA}}_{\text{locality}}$ and $\mathcal{S}^{\text{QA}}_{\text{mhop}}$ are used for evaluation, while $\mathcal{S}^{\text{QA}}_{\text{changed}}$ constitutes the respective fact-based training data.
(6) Generation of Personalized and Rephrased Question-Answer Edit Pairs for Generalization Studies. While $\mathcal{S}^{\text{QA}}_{\text{locality}}$ and $\mathcal{S}^{\text{QA}}_{\text{mhop}}$ allow us to probe locality and reasoning generalization of edits, we also provide $\mathcal{S}^{\text{QA}}_{\text{rephrase}}$ and $\mathcal{S}^{\text{QA}}_{\text{personas}}$ containing rephrased variants of $\mathcal{S}^{\text{QA}}_{\text{changed}}$ to check generalization of edits beyond memorization. In particular, we utilize GPT-4o-mini to provide additional question-answer variations of $\mathcal{S}^{\text{QA}}_{\text{changed}}$ by enforcing the language model to either directly rephrase $\mathcal{S}^{\text{QA}}_{\text{changed}}$ to produce $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$ or directly generate $\mathcal{S}^{\text{QA}}_{\text{personas}}$ by mimicking various different “personas” (e.g. “detective”, “pirate”, “philosopher” defined by a detailed description, see supplementary). For both $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$ and $\mathcal{S}^{\text{QA}}_{\text{personas}}$, post-hoc filtering is applied to ensure that subject, object and relation remain correctly encoded in both.
(7) Final Filtering Stage. A final filtering process ensures the quality of the generated question-answer pairs for both training and evaluation sets by verifying that the generated questions contain the respective subject and relation while not giving away the answer and answers incorporate the right object. After this stage, the final $\mathcal{S}^{\text{QA}}_{\text{changed}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$, $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$ an $\mathcal{S}^{\text{QA}}_{\text{personas}}$ are incorporated as a new knowledge edit batch into WikiBigEdit.
3.2 Analysis and Exploration
Timestep Date Range Samples Unsolved
T1 2024/02/01 - 2024/02/20 26,922 80%
T2 2024/02/20 - 2024/03/01 29,835 84%
T3 2024/03/01 - 2024/03/20 54,504 85%
T4 2024/03/20 - 2024/04/01 43,443 85%
T5 2024/04/01 - 2024/05/01 121,116 81%
T6 2024/05/01 - 2024/06/01 101,728 82%
T7 2024/06/01 - 2024/06/20 69,403 82%
T8 2024/06/20 - 2024/07/01 55,431 82%
Total 502,382 82%
Table 2: Description of WikiBigEdit timesteps. Sample counts reflect number of question-answer pairs associated with a given time-interval; alongside information highlighting the percentage of unsolved edits across studied LLMs; highlighting the large number of required real-world factual updates across LLMs.
Figure 2: Analysis and Exploration of Real-World Factual Updates in WikiBigEdit. (a) Distribution of time-stamps associated with factoids where temporal information was available. For real-world updates, facts cover both historical, but especially more recent context. (b) Average accuracy of five LLMs evaluated across facts broken down by timestamps. We find that contrasted against facts clearly preceding knowledge cutoffs (pre-2023), there is a notable drop in average capabilities, highlighting the need to both address edits on pretraining context, but especially the inclusion of recent facts or corrections. (c) Distribution of edits extracted for each timestep and their corresponding dates. While the majority of edits align with their respective timestep, edits in larger parts also incorporate changes of information associated with primarily “older” time intervals. (d-g) Real-world edits cover both highly specific facts (measured on subject and object counts in pretraining corpora) as well as general, common knowledge modifications. (h-i) Especially on highly specific edits as defined by pretraining concept frequency does performance go down.
In this section, we first summarize various benchmark characteristics before leveraging the first instantiation of our large-scale WikiBigEdit to study the structure and properties of real-world knowledge edits derived from Wikidata knowledge graphs. Table 2 provides an overview of the eight time-intervals that constitute WikiBigEdit. The factual updates are extracted on a roughly bi-monthly frequency from Wikidata snapshots spanning February to July 2024; with the availability of Wikidata snapshots determining size and extend of factual updates associated with each timestep.
To ensure high data quality, we filter out approximately 76% of initial changes, removing cyclic or overly long entities (11%), non-entity-to-entity relations (50%), and nondeterministic triplets (15%). The filtered updates range from 27k to 69k per step, with monthly factual changes exceeding 100k. This variability reflects the dynamic nature of knowledge updates and practical constraints when aiming to ensure that models remain up-to-date. Overall, this version contains 502,382 question-answer pairs across 6.9M tokens.
Impact of Language Model Choice on QA Generation The quality of generated question-answer pairs in WikiBigEdit depends in part on the underlying language model used during generation. Before we conduct further experiments, we begin by examining the sensitivity of the generation process to different LLMs by comparing outputs from GPT-3.5-turbo, GPT-4o-mini, and GPT-4o. For both the Update and Locality sets, which are directly derived from structured subject–relation–object triplets, model choice has negligible impact. Even the less capable GPT-3.5-turbo consistently produced well-formed, semantically correct questions. As shown in Table 3, the outputs across all models for such simple factual prompts are stylistically comparable and accurate. In contrast, the Multi-hop and Persona sets were more sensitive to model choice, as they require compositional reasoning and stylistic transformation. For instance, GPT-3.5-turbo often struggled to combine facts effectively in multi-hop settings, producing incomplete or ambiguous questions. GPT-4o-mini and GPT-4o, in comparison, reliably produced complete and coherent outputs. These observations confirm our selection for GPT-4o-mini to generate Multi-hop and Persona sets due to its strong balance between generation quality and computational efficiency.
Update Triplet Lea County Regional Airport - state of use - in use
GPT-3.5 What is the current state of use for Lea County Regional Airport?
GPT-4o-mini What is the current state of use of Lea County Regional Airport?
GPT-4o What is the current state of use of Lea County Regional Airport?
Mhop Tuple 2004 VS75 - discoverer - Marc Buie - gender - male
GPT-3.5 Who is the discoverer or inventor of 2004 VS75?
GPT-4o-mini What is the sex or gender of the discoverer of 2004 VS75?
GPT-4o What is the gender of the discoverer or inventor of 2004 VS75?
Table 3: Comparison of generated questions across different language models for a simple factual update (top) and a multi-hop reasoning example (bottom). While all models perform well on the single-hop case, GPT-3.5 struggles with compositional reasoning, often producing incomplete multi-hop questions. GPT-4o-mini and GPT-4o generate more complete and semantically accurate queries.
Understanding the temporal resolution of real-world factual updates. To assess the temporal resolution of knowledge edits in our benchmark, we analyze timestamps for factual triplets where date information is available (10,409 question-answer pairs). As shown in Figure 2 (a), updates strongly focus on recent events, though many also modify facts predating the edit dates. Figure 2 (c) further reveals that extracted updates are not strictly confined to their respective timesteps, underscoring the dual challenge of integrating both new knowledge and modifying existing facts from pretraining corpora. When stratified by knowledge cutoff (pre-2023 vs. 2023/2024), all tested LLMs perform worse on post-cutoff facts (Cheng et al., 2024) (b), highlighting the critical need for effective integration of new knowledge.
Specificity of factual updates. We assess the specificity of knowledge edits by distinguishing well-known from highly specialized facts, using two measures: (1) entity frequency from the Infinigram API (Liu et al., 2024a), and (2) Wikipedia view counts (Jan 2023–June 2024) (Mallen et al., 2023). As shown in Figure 2 (d–g), WikiBigEdit includes both general (e.g., Which continent is Krasnodar a part of?) and highly specific questions (e.g., Who was the doctoral advisor of Omar Khayyám?). To evaluate how specificity affects performance, we test five LLMs (Llama-2-7b, Llama-3-8b, Mistral-7b, Gemma-7b, xGen-7b) and correlate their accuracy. As shown in (h), lower entity frequency correlates with reduced performance, consistent with Udandarao et al. (2024); Cheng et al. (2024), while view counts (i) show less consistent trends. Specificity appears primarily driven by the object rather than the subject. See supplementary for details.
WikiBigEdit - Analysis Summary When evaluating LLMs on a large number of real-world factual updates, performance is at best mediocre, but especially low for highly specific and recent facts. For LLMs to serve as accurate and timely knowledge bases, the ability to edit knowledge across historic, recent or specific context at scale is thus crucial.
4 Understanding the Limits of Lifelong Knowledge Editing in LLMs
Leveraging WikiBigEdit, we evaluate the ability of lifelong knowledge editing techniques to incorporate large-scale factual updates. Section 4.1 presents a comprehensive application of various lifelong editing approaches and compares them to alternative methods like continual finetuning and retrieval augmentation. Section 4.2 analyzes the specific limitations of knowledge editing techniques, with Section 4.3 and Section 4.4 providing detailed evaluations of retrieval augmentation and continual finetuning.
Experimental Details. All experiments are performed on a compute cluster equipped with Nvidia A100 and H100 GPUs, leveraging PyTorch (Paszke et al., 2019) and building on the EasyEdit codebase (Zhang et al., 2024). All our evaluations are conducted on five state-of-the-art LLMs: Llama-2-7b (et al., 2023) Llama-3-8b (et al., 2024a), Mistral-7b (Jiang et al., 2023), Gemma-7b (Team et al., 2024a) and xGen-7b (Nijkamp et al., 2023). To representatively study knowledge editing, we leverage the following approaches: ROME (Meng et al., 2023), R-ROME (Gupta et al., 2024a), MEMIT (Meng et al., 2022), and WISE (Wang et al., 2024a). For our retrieval-augmented generation experiments, we store update question-answer pairs in memory alongside associated key embeddings generated by embedding questions with the all-mpnet-base-v2 embedding model (Song et al., 2020), the best general sentence-embedding model provided in the SBERT repository (Reimers & Gurevych, 2019). During inference, the top-2 most similar entries are retrieved from memory efficiently using the Annoy solver (Bernhardsson, 2018). This setup ensures efficient retrieval and supports multi-hop questions with minimal computational overhead. For our continual finetuning experiments, we explore variations of low-rank adaptation using LoRA (Hu et al., 2021) alongside simple interpolation-based model-merging (Wortsman et al., 2022; Stojanovski et al., 2022; Marouf et al., 2024) which has shown to perform competitively even when compared to more involved merging techniques (Roth et al., 2024; Dziadzio et al., 2024). Adapters are trained for 10 epochs on each timestep. Additional details are provided in the corresponding subsections and the supplementary.
4.1 Lifelong Knowledge Editing on WikiBigEdit
Figure 3: Large-Scale Lifelong Knowledge Editing on WikiBigEdit. Comparison of knowledge editing techniques and other standards for model modification (retrieval-augmented generation, RAG, and continual finetuning using LoRA + merging). Results reveal that RAG vastly outperforms specialized knowledge editing techniques (at higher inference cost). At equivalent inference, simple continual finetuning consistently improves on advanced editing techniques at scale.
To begin, Figure 3 presents a high-level comparison of the average accuracy across all eight timesteps $U_{1},...,U_{8}]$ for all five LLMs listed above. After incorporating factual changes of a timestep, evaluation is conducted across all five evaluation axes spanned by respective evaluation data $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$, $\mathcal{S}^{\text{QA}}_{\text{personas}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$, and a performance evaluation on the actual edit data $\mathcal{S}^{\text{QA}}_{\text{update}}$. Results across all models and timesteps are averaged and visualized.
RAG (green line) proves to be the most effective approach, significantly outperforming knowledge editing techniques across evaluation axes, nearly tripling accuracy on training edits. This result is expected, as memorizing previously seen edits makes solving the task straightforward. Memorization also boosts performance on generalization and locality metrics, commonly used in knowledge editing literature. However, compositional reasoning on multihop questions sees only slight improvements, as expected.
RAG naturally trades performance increases for longer inference times. However, with effective approximate solvers and high-quality sentence embeddings, we find that inference time at most doubles after processing all of WikiBigEdit (c.f. LABEL:fig:forward_pass_time). Even so, knowledge editing techniques like WISE and MEMIT, specifically designed to factually update LLMs, cannot match simple low-rank adapter-based finetuning (LoRA-FT) at scale. Moreover, with interpolation-based model merging (LoRA-Merge), commonly used in training and finetuning LLMs (Yang et al., 2024a; Roth et al., 2024; Team et al., 2024a, b; Dziadzio et al., 2024), consistent gains are observed across all evaluation axes. These results strongly suggest that at scale, standard knowledge editing techniques fall short even when compared to straightforward continual finetuning approaches commonly employed for general continual pretraining (Ibrahim et al., 2024; Roth et al., 2024).
Key Takeaway Updated via WikiBigEdit, RAG significantly outperforms all approaches at double the inference time; though expectedly struggles with multi-hop reasoning. Given a fixed inference budget, we find that at scale, knowledge editing approaches evaluated across all standard editing evaluation axes are actually matched or outperformed by simple continual finetuning, especially using merging.
The following sections provide a more detailed breakdown of specific failure scenarios and drivers.
4.2 The Limits of Knowledge Editing at Scale
Figure 4: Performance Decay of Knowledge Editing Techniques. Update accuracy over first 10K edits for knowledge editing methods compared to simple baselines that show higher performance (LoRA, RAG). Local modification methods collapse early (MEMIT, ROME, R-ROME), and even those specifically designed for large volumes of edits (WISE) converge to pre-edit accuracy.
This section provides an in-depth analysis of knowledge editing applied to long and realistic update sequences, with results shown in Figure 4. Using Llama-2-7b as the underlying LLM, we present the first 10K update steps from the initial timestep in WikiBigEdit for clarity. A complete version with all models is available in the supplementary.
The accuracy of ROME, R-ROME, and MEMIT - local knowledge editing techniques that locally modify specific model weights - rapidly degrades within the first few hundred updates, leading to model collapse. This failure on large-scale, sequential knowledge editing aligns with recent findings (Gupta et al., 2024b; Yang et al., 2024b) and highlights difficulties for lifelong knowledge editing at scale. Even without direct model edits, methods like WISE - designed for larger update volumes by storing adapted weight matrices for retrieval during inference - show limited benefits. While WISE avoids catastrophic failure and integrates new facts over longer update sequences, its performance steadily declines within the first 10K updates, ultimately converging to pre-update knowledge levels. This suggests that by avoiding direct model modifications, WISE struggles to retain updates over time, making it unsuitable for large-scale sequential integration, especially compared to retrieval augmentation and simple continual fine-tuning, which incur no additional inference cost.
Key Takeaway Knowledge editing is reliant on local weight modification and collapses quickly. Even methods designed for lifelong editing converge to pre-update performance early. So large-scale lifelong editing requires significantly improving existing methods.
4.3 Retrieval Augmentation for WikiBigEdit
Figure 5: Retrieval-Augmented Generation for Lifelong Knowledge Editing. (Top) Mean batch accuracy for all evaluations: update, rephrase, personas, multi-hop and locality across all models. (Bottom) Accuracy changes for preceding timesteps after training on a given timestep, highlighting forgetting symptoms on generalization tasks, and marginal benefits at best for reasoning over edits.
As shown in Figure 3, retrieval-augmented generation (RAG) demonstrates significant improvements over direct knowledge editing techniques. While expected for performance on training edits, the increased performance across all evaluation axes, including generality, is somewhat surprising, as prior retrieval approaches like GRACE (Hartvigsen et al., 2023) have shown limited generalization capabilities (Wang et al., 2024a). However, on realistic edits and at scale, we find that efficiently implemented retrieval-augmentation outperforms any editing method (see Section 4 and supplementary for details).
Looking at the detailed results of the RAG baseline on WikiBigEdit in Figure 5, the improvements across all evaluation axes (except multi-hop reasoning) are significant across 500K+ updates. The top row shows the average accuracy on the current timestep, with bold lines representing average performance across five models (light background lines). The bottom row illustrates average performance across all timesteps after training on the current timestep’s edits. As noted previously, RAG achieved near-perfect accuracy on $\mathcal{S}^{\text{QA}}_{\text{update}}$ by simply retrieving the exact stored questions from memory; exhibiting no sign of “forgetting” on trained edits (c.f. bottom row, leftmost). Additionally, edits remain highly localized since no weight updates are performed. Interestingly, while RAG also maintains high average performance on generalization-based evaluation axes (“rephrasing” and “personas”), achieving average accuracies of 78.64% and 66.16%, respectively, RAG still exhibits symptoms akin to catastrophic forgetting encountered in continual learning and knowledge editing (Kirkpatrick et al., 2017; Zenke et al., 2017; Wang et al., 2024a). Here, performance on older timestep data decays with every new large update batch (bottom, second and third), as the retrieval memory becomes more and more obfuscated with context irrelevant for data from particular timesteps.
Finally, while RAG outperforms knowledge editing approaches on standard measures (locality, generalization, and overall edit performance), its multi-hop reasoning accuracy remains near pre-update levels (second from right in Figure 5). Figure 6 reveals the bottleneck: retrieval accuracy is low - likely due to the absence of the middle subject in the multi-hop question - with correct retrieval rates below 50% for the first hop and even lower for the second. Only 10% of evaluation samples retrieve both facts correctly. Even when both hops are retrieved, answer accuracy remains near single-hop levels (i.e. solving multi-hop questions using only information from just the first hop), as models struggle to combine retrieved information without dedicated training for reasoning tasks. While advanced methods like HippoRAG (Gutiérrez et al., 2025) may help, they typically require longer contexts or higher compute; we leave their integration to future work.
Key Takeaway Simple RAG outperforms much more involved knowledge editing techniques across all editing evaluation axes at larger scale and real-world edits. However, performance on evaluation over older edits still decays, and reasoning over edited facts shows marginal benefits at best.
Figure 6: Multi-hop RAG. (Left) Retrieval accuracy for individual and both hops, highlighting significant challenges in retrieving both relevant facts simultaneously. (Right) Answer accuracy conditioned on correct retrieval, showing that even when both hops are retrieved, models struggle to combine the information.
4.4 Continual Finetuning
Figure 7: Continual Finetuning for Knowledge Editing. (Top) Mean batch accuracy for update, rephrase, personas, mhop, and locality sets for continual LoRA finetuning (LoRA-FT) with model merging (LoRA-Merge). (Bottom) Accuracy differences for individual timesteps. LoRA-FT degrades over time, while merging ensures consistent editability across all WikiBigEdit.
To conclude, we investigate continual finetuning using low-rank adaptation (LoRA) as a simple yet effective method to incorporate large volumes of new facts. While recent works (e.g., Thede et al. (2024); Roth et al. (2024)) have demonstrated the benefits of continual low-rank finetuning for general continual learning tasks, we show that these advantages extend to specialized knowledge editing for LLMs. As shown in Figure 3, simple non-specialized finetuning over all WikiBigEdit updates outperforms specialized knowledge editing techniques across all standard evaluation axes, with gains further amplified by interpolation-based model merging. These findings emphasize the need to revisit design choices in knowledge editing, evaluate on larger, realistic benchmarks, and improve baseline comparisons.
Finally, we provide a detailed perspective on lifelong knowledge editing using continual finetuning in Figure 7 (rank-4 LoRA; additional results in the supplementary [^5] High-rank configurations exhibit significant instability, including catastrophic failure on Mistral and poor performance across other models.). As shown in the top row, performance remains competitive in early timesteps but steadily declines as updates accumulate, eventually converging to or even falling below pre-update accuracy, particularly for locality and multi-hop reasoning. This aligns with expectations, as catastrophic forgetting persists even in low-rank adaptation (Thede et al., 2024; Roth et al., 2024). However, for the first 100K updates, continual finetuning shows notable success in integrating edits. These results highlight its viability for mid-term knowledge integration but also its limitations at scale.
Interestingly however, as we couple continual LoRA finetuning with simple model merging (after training of each timestep, current adapter weights are simply merged into preceding adapter weights using an interpolation weight of 0.25), we achieve significantly improved editing performance across WikiBigEdit. While LoRA-Merge initially shows lower performance on shorter update sequences, it demonstrates superior stability as updates accumulate; and begins to notably outperform simple continual finetuning (and thereby also any specialized knowledge editing technique) after approximately 100K edits on all evaluation sets, maintaining consistent performance without significant degradation. Importantly, LoRA-Merge shows at most marginal evidence of catastrophic forgetting, and in parts even exhibits positive backward transfer (where training on a current timestep actually improves performance on all edits seen previously) - see Figure 7 bottom row. These results reinforce the effectiveness of model merging for general continual finetuning and pretraining tasks (Stojanovski et al., 2022; Thede et al., 2024; Roth et al., 2024), highlighting its promise for large-scale knowledge editing.
Key Takeaway Generic adapter-based finetuning matches specialized knowledge editing for mid-level knowledge editing; though exhibits catastrophic forgetting at scale. However, equipped with simple model merging, we find consistent inclusion of new facts across all of WikiBigEdit. This suggests better baseline comparisons and studies at scale for the development of specialized knowledge editing techniques.
5 Conclusion
This paper explores lifelong knowledge editing for large language models at a practical scale. We introduce WikiBigEdit, a benchmark with an automated data acquisition pipeline that extracts real-world factual updates from Wikidata, enabling lifelong expansion. In its first version, WikiBigEdit contains over 500K QA pairs for large-scale lifelong knowledge editing, complemented by a comprehensive evaluation suite covering standard knowledge editing criteria, reasoning, and generalization.
Leveraging WikiBigEdit, we conduct an extensive experimental study in which we identify key limitations of existing knowledge editing approaches such as ROME, MEMIT, and WISE, which struggle with scalability and stability in large-scale update settings. Our results also demonstrate the efficacy of retrieval-augmented generation (RAG) when no restriction on inference time is given and highlight continual finetuning coupled with model merging as a competitive alternative at an equivalent inference cost.
We believe that WikiBigEdit provides an important reference to understand the transfer of knowledge editing methods to real-world scale and facts, as well as suggestions towards more complete baseline studies during development of such specialized techniques.
Acknowledgements
LT and KR thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. KR also thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. We are grateful for support by the Carl Zeiss Foundation, project ”Certification and Foundations of Safe Machine Learning Systems in Healthcare”. This work was partially funded by the ERC (853489 - DEXIM) and the Alfried Krupp von Bohlen und Halbach Foundation, which we thank for their generous support.
Impact Statement
Maintaining factual accuracy in large language models (LLMs) is critical, as these models are increasingly relied upon in domains where misinformation or outdated knowledge can have serious societal consequences. By focusing on the lifelong integration of factual knowledge, this paper directly addresses the challenge of ensuring that LLMs remain up-to-date, reliable, and accurate over time. The proposed WikiBigEdit benchmark provides a valuable tool for designing and evaluating methods that enhance factuality, offering a scalable and realistic framework grounded in real-world knowledge evolution. By advancing techniques for continual factual updates, this work contributes to reducing the risks of outdated or incorrect information, promoting trust and accountability in the use of LLMs for applications in education, healthcare, journalism, and beyond.
Appendix A Benchmark Generation
This section provides additional details about the benchmark generation process, focusing on the templates used to create question-answer pairs from the extracted factual updates. Each subset of the benchmark — update, locality, rephrase, personas, and multi-hop — is generated using carefully designed prompts tailored to their specific evaluation objectives. Below, we describe the prompt templates and methodologies for generating these subsets.
A.1 Update and Locality Set
The update set serves as the foundation of our benchmark, containing questions derived directly from the extracted factual updates. The locality set, on the other hand, evaluates the robustness of knowledge updates by ensuring that unrelated but similar facts remain unaffected.
We employ the following prompt template to generate both the update and locality sets. The prompt includes a general task description, detailed instructions, and few-shot examples to guide the model in generating high-quality question-answer pairs:
Prompt Template Update/Locality Given a fact triplet in the format (subject, relation, object) and a description of the relation, generate a question-answer pair where the answer is the object of the triplet. Ensure that the question explicitly references the relation and is structured to lead uniquely to the chosen answer, which must be the exact text of the object. Use the description of the relation to improve the specificity and clarity of the question. Make the question specific enough to ensure that there is only one possible correct answer, eliminating any ambiguity.
Instructions for Generation: 1. Construct the Question: Ensure the question is directly tied to the relation and points exclusively to the selected answer. Include additional context from the subject to make the question more specific and unambiguous. Use the description of the relation to enhance the quality of the question. Use only the exact text of the object in the answer. 2. Format: Provide the question-answer pair in the following format: Q: A: Examples: Fact Triplet: Turnberry Lighthouse, color, white
Relation Description: The color of the subject.
Q: What is the color of Turnberry Lighthouse?
A: white
Fact Triplet: Folsom Library, part of, Rensselaer Libraries
Relation Description: part of this subject; inverse property of “part of” (P361). See also “has parts of the class” (P2670).
Q: What institution does Folsom Library belong to?
A: Rensselaer Libraries
Fact Triplet: Volt Europa, subsidiary, Volt Netherlands
Relation Description: A company controlled by another company.
Q: What is the name of the subsidiary of Volt Europa in the Netherlands?
A: Volt Netherlands
Fact Triplet: Zonia Baber, educated at, Chicago State University
Relation Description: The institution where the subject received their education.
Q: At which university did Zonia Baber receive her education?
A: Chicago State University
Task: According to the instructions and examples, provide a QA pair for the following fact triplet:
Fact Triplet: {}, {}, {}
Relation Description: {}
A.2 Rephrase Set
The rephrase set is designed to test the generalization of updated knowledge by reformulating the questions from the update set. The reformulated questions retain the same factual context but vary in phrasing, introducing linguistic diversity to the benchmark.
The following prompt is used to generate the rephrased questions:
Prompt Template Rephrase You are given a question-answer pair. Reformulate the question such that its content remains identical and the given answer is still the only accurate answer.
Instructions for Generation: 1. Question Reformulation: Restructure the question while maintaining the same content and context. Ensure that the answer provided in the original question remains the only correct answer. 2. Response Format: Provide the reformulated question in the following format: Reformulated Question: Examples: Original Question: What is the capital of the United States?
Answer: Washington, D.C.
Reformulated Question: What city serves as the capital of the United States?
Original Question: Who painted the Mona Lisa?
Answer: Leonardo da Vinci
Reformulated Question: By whom was the Mona Lisa painted?
Original Question: When was the Declaration of Independence signed?
Answer: July 4, 1776
Reformulated Question: On what date was the Declaration of Independence signed?
Task: Original Question: {} Answer: {}
A.3 Persona Set
The persona set adds a layer of complexity by requiring the model to generate questions in the style of distinct personas. Each question from the update set is rewritten based on the characteristics of a randomly selected persona.
Table 4: Results of WISE on three language models (Llama-2-7B, Llama-3-8B, and Mistral-7B) for the first timestep of the WikiBigEdit benchmark. Each subplot shows mean batch accuracy for the update, rephrase, personas, mhop, and locality sets over the number of updates. Post-update performance (blue) converges to pre-update levels (gray) across all metrics, highlighting WISE’s limitations in sustaining accuracy after large-scale updates.
Persona Description
Detective You are a world-weary detective narrating your investigations. Your tone is gritty, mysterious, and evocative of noir fiction. Reformulate the following questions as if you’re puzzling out clues in a case. Keep the core meaning intact but phrase it in your distinct detective style.
Casual You are a casual, friendly conversationalist with a relaxed and approachable tone. You’re curious and engaging, as if you’re chatting with a friend over coffee or in a group discussion. Reformulate the following questions to sound natural and conversational while keeping the core meaning intact.
Pirate You are a swashbuckling pirate with a flair for colorful language and sea-themed metaphors. Your tone is bold, rough, and full of pirate lingo, as if you’re recounting tales from the high seas. Reformulate the following questions to sound like they’re being asked by a true pirate. Keep the core meaning intact but add a piratical twist.
Philosopher You are an ancient philosopher, deeply contemplative and wise, always pondering the greater truths of existence. Your tone is reflective, profound, and formal, as if you are crafting a dialogue or treatise on the nature of knowledge and events. Reformulate the following questions as if they are inquiries befitting philosophical discourse. Keep the core meaning intact but phrase it in your distinct, thoughtful style.
Caveman You are a caveman, speaking in simple and primitive language, with short sentences and limited vocabulary. Your tone is direct, straightforward, and reflects the early stages of human communication. Reformulate the following questions in a way that matches your basic and minimalistic speech style, while keeping the core meaning intact.